# Adaptive Patching for Tensor Train Computations

### v1

@@ 5.1 Matsubara axis
We consider a 2D square lattice with the bare Green’s function presented in (15), $G({\bm{k}},i\nu_{n})=(i\nu_{n}+\mu-\varepsilon_{\bm{k}})^{-1}$, which we QTT-compress jointly over momentum and Matsubara frequency, as the starting point for evaluating $\chi_{0}(\bm{q},i\omega)$. The complete workflow is depicted in the flowchart of Figure 12. The momentum dependence of the input Green’s function and the resulting bare susceptibility at the first and zeroth Matsubara frequency respectively, as well as the distribution of patches for the Fourier transformed Green’s function $\hat{G}(\bm{r},\bar{\tau}=\beta)$ and bond dimension per patch are illustrated in Fig. 13. The sharp arc-like structures present in $G({\bm{k}},i\pi/\beta)$ at $\beta=100$ (cf. Fig. 13 (a)) reflect the shape of the Fermi surface in the doped 2D Hubbard model, which is almost incompressible (cf. large bond dimensions in Fig. 13 (d)). The sharpness of the features, and hence the bond dimension, increases with inverse temperature $\beta$; see also Fig. 3 where the broadening $\delta$ plays the same role as temperature here. After patching, the bond dimension per patch is significantly lower. Remarkably, this reduces the runtime of the bubble calculation by an order of magnitude, as shown in Fig. 14(b). For $\mu=0$, the Fermi surface is a square (not shown) leading to slightly lower bond dimensions and shorter runtimes both for the patched and unpatched implementations (cf. Fig. 14(a)), compared to the $\mu=1$ case.
@@
Figure 13: Heatmap overview of the bare-susceptibility workflow with $\mathcal{R}=13$, $\beta=100$, $\mu=1$ and tolerance $\tau=10^{-5}$. (a) Absolute value of the Green’s function $G({\bm{k}},i\pi/\beta)$ at the first positive fermionic Matsubara frequency. (b) Bare susceptibility $\chi_{0}(\bm{q},0)$ (c) Adaptive patching pattern for $\hat{G}(\bm{r},\beta)$. (d) From left to right, bond dimension profile of TCI approximation of $G({\bm{k}},i\pi/\beta)$ (cf. (a)), bond dimension profile of $\chi_{0}(\bm{q},0)$ QTT (cf. (b)) and bond dimension of each patch for the pQTCI approximation of $\hat{G}(\bm{r},\beta)$ (cf. (c)).
Figure 14: CPU time for the patched element-wise product compared with the monolithic contraction as a function of tolerance $\tau$ and at different inverse temperatures $\beta$ for $\mu=0$ (a) and $\mu=1$ (b). The benchmark was performed on as single Intel® Xeon® E5-2680 v4 @ 2.40 Ghz.

### v2

@@ 5.1 Matsubara axis
We consider a 2D square lattice with the bare Green’s function presented in (15), $G({\bm{k}},i\nu_{n})=(i\nu_{n}+\mu-\varepsilon_{\bm{k}})^{-1}$, which we QTT-compress jointly over momentum and Matsubara frequency, as the starting point for evaluating $\chi_{0}(\bm{q},i\omega_{n})$. The complete workflow is depicted in the flowchart of Figure 12. The index ordering for the QTTs is chosen as follows: in momentum space, the $k_{x}$ and $k_{y}$ quantics legs are interleaved and the Matsubara frequency legs are appended at the end; in real space, the imaginary time quantics legs are placed at the beginning of the TT and the $x$ and $y$ legs are interleaved thereafter. Representative frequency and imaginary-time slices of $G({\bm{k}},i\nu_{n})$, $\hat{G}(\bm{r},\bar{\tau})$ and $\chi_{0}(\bm{q},i\omega_{n})$ are shown as heatmaps in Fig. 13(a),(c),(e), with the bond dimension profiles of the corresponding full objects shown in Fig. 13(b),(d),(f). The sharp arc-like structures present in $G({\bm{k}},i\pi/\beta)$ at $\beta=100$ (cf. Fig. 13 (a)) reflect the shape of the Fermi surface in the doped 2D Hubbard model, which is almost incompressible (cf. large bond dimensions in Fig. 13 (b)). The sharpness of the features, and hence the bond dimension, increases with inverse temperature $\beta$; see also Fig. 3 where the broadening $\delta$ plays the same role as temperature here. After patching — which is applied exclusively to the real space quantics indices of $\hat{G}(\bm{r},\bar{\tau})$ (i.e. starting from the $\mathcal{R}$th leg), as evidenced by the flat intermediate regions in the bond dimension profile in Fig. 13(d) — the bond dimension per patch is significantly lower. Remarkably, this reduces the runtime of the bubble calculation by an order of magnitude relative to the unpatched computation, as shown in Fig. 14(b). For $\mu=0$, the Fermi surface is a square (not shown) leading to slightly lower bond dimensions and shorter runtimes both for the patched and unpatched implementations (cf. Fig. 14(a)), compared to the $\mu=1$ case.
@@
Figure 13: Heatmap and bond dimensions overview of the bare-susceptibility workflow with $\mathcal{R}=13$, $\beta=100$, $\mu=1$ and tolerance $\tau=10^{-5}$. Left column: representative slices of each object. Right column: bond dimension profiles of the corresponding full objects. (a) Absolute value of $G({\bm{k}},i\pi/\beta)$ at the first positive fermionic Matsubara frequency. (b) Bond dimension profile of the TCI approximation of $G({\bm{k}},i\nu_{n})$. (c) $\hat{G}(\bm{r},\beta)$ in real space at imaginary time $\bar{\tau}=\beta$. (d) Bond dimension profiles of $\hat{G}(\bm{r},\bar{\tau})$ for the unpatched and patched (pQTCI) approximations; the bond-dimension cap is set to $\chi_{p}=181$. (e) Bare susceptibility $\chi_{0}(\bm{q},0)$ at the zeroth bosonic Matsubara frequency. (f) Bond dimension profile of the QTT approximation of $\chi_{0}(\bm{q},i\omega_{n})$.
Figure 14: CPU time for the patched element-wise product compared with the monolithic contraction as a function of tolerance $\tau$ and at different inverse temperatures $\beta$ for $\mu=0$ (a) and $\mu=1$ (b). For each set of parameters the optimal bond-dimension caps $\chi_{\text{p}}$ are chosen (cf. Figs. 4, 5). The benchmark was performed on a single Intel® Xeon® E5-2680 v4 @ 2.40 Ghz.